# Accuracy of the Yee FDTD Scheme for Normal Incidence of Plane Waves on Dielectric and Magnetic Interfaces

### v1

@@ 6.2 Magnetic/magnetic interface
Thus, the expressions (66) and (67) prove that the equality (65) is the first equation in (62). Its substitution into the boundary condition (47) allows us to express the coefficient $\widetilde{r}$ in the form of the second equation (62), which completes the proof. ∎
@@ 7.1 Peculiarities of the FDTD Fresnel coefficients
Moreover, Fig.  5 contains information about the two types interfaces of dielectrics and magnetics: the dashed lines represent the dependences on $N_{\lambda}$ of the FDTD coefficients $\widetilde{r}$ and $\widetilde{t}$ when $\varepsilon_{1}<\varepsilon_{2}$ (red curves for dielectrics interface) and $\mu_{1}<\mu_{2}$ (blue lines in the magnetics example), while the point-dashed lines correspond to the opposite situations.
@@ 7.3 Possible topics for further research
The asymptotic behavior of the Fresnel coefficients, as well as the reflection and transmission coefficients in the $N_{\lambda}\rightarrow\infty$ region for the FDTD interface model shown in Fig. 11 also tends to be their exact analogs to the same extent as for the models shown in Fig. 3 and 4. Moreover, numeric error remains $O(\Delta_{x}^{2})$ and Corollaries 1, 2 and 4 qualitatively persist when impedance contrast dominates.
@@ 8 Conclusion
§4 contains all the necessary general information about the FDTD method, including the dispersion relation for the Yee grid, which is used later to formulate the finite-difference analogue of the same problem. In particular, subsection §4.1 provides the FDTD analogues of the Faraday (19) and Ampère-Maxwell (21) laws equations. Subsection §4.2 contains a description of the process of generating the incident wave field in the dynamic mode in the TFSF formalism, expressed in the form of equations (24) and (25). Expressions for FDTD analogs of the incident (26), reflected (27) and transmitted (28) waves, as well as their main features, are formulated further in §4.3. Finally, in §4.4, the dispersion relation equation for the Yee grid presented in Lemma 1 is given.
The main research part of the work is concentrated further in §5, §6 and §7. In §5 two basic FDTD models are defined for the interfaces between dielectrics and magnetics, within the scope of which two theorems on the corresponding boundary conditions are formulated and proven. Subsection §5.1 contains necessary definitions and lemmas, which are used further to obtain the main results of this section. In §5.2, it is proven that equations (40) and (41) are FDTD analogues of the boundary conditions relating the tangential components of the electromagnetic field of a plane monochromatic wave at the interface of two dielectrics. The validity of equations (47) and (48), which are FDTD analogues of the boundary conditions at the interface of two magnetics, is proven in §5.3.
@@ Acknowledgments
The authors express their sincere gratitude to John B. Schneider, Associate Profesor of Washington State University for useful discussion of FDTD disperion effects and all anonymous reviewers for their valuable suggestions that improved this work.

### v2

@@ 6.2 Magnetic/magnetic interface
Thus, the expressions (66) and (67) prove that the equality (65) is the first equation in (62). Its substitution into the boundary condition (47) allows us to express the coefficient $\widetilde{r}$ in the form of the second equation in (62), which completes the proof. ∎
@@ 7.1 Peculiarities of the FDTD Fresnel coefficients
Moreover, Fig. 5 contains information about the two types interfaces of dielectrics and magnetics: the dashed lines represent the dependences on $N_{\lambda}$ of the FDTD coefficients $\widetilde{r}$ and $\widetilde{t}$ when $\varepsilon_{1}<\varepsilon_{2}$ (red curves for dielectrics interface) and $\mu_{1}<\mu_{2}$ (blue lines in the magnetics example), while the point-dashed lines correspond to the opposite situations.
@@ 7.3 Possible topics for further research
The asymptotic behavior of the Fresnel coefficients, as well as the reflection and transmission coefficients in the $N_{\lambda}\rightarrow\infty$ region for the FDTD interface model shown in Fig. 11 also tends to be their exact analogs to the same extent as for the models shown in Fig. 3 and 4. Moreover, numeric error remains the same and Corollaries 1, 2 and 4 qualitatively persist when impedance contrast dominates.
@@ 8 Conclusion
§4 contains all the necessary general information about the FDTD method, including the dispersion relation for the Yee grid, which is used later to formulate the finite-difference analogue of the same problem. In particular, subsection §4.1 provides the FDTD analogues of the Faraday (20) and Ampère-Maxwell (22) laws equations. Subsection §4.2 contains a description of the process of generating the incident wave field in the dynamic mode in the TFSF formalism, expressed in the form of (24) and (25). Expressions for FDTD analogs of the incident (26), reflected (27) and transmitted (28) waves, as well as their main features, are formulated further in §4.3. Finally, in §4.4, the dispersion relation equation for the Yee grid presented in Lemma 1 is given.
The main research part of the work is concentrated further in §5, §6 and §7. In §5 two basic FDTD models are defined for the interfaces between dielectrics and magnetics, within the scope of which two theorems on the corresponding boundary conditions are formulated and proven. Subsection §5.1 contains necessary definitions and lemmas, which are used further to obtain the main results of this section. In §5.2, it is proven that (40) and (41) are FDTD analogues of the boundary conditions relating the tangential components of the electromagnetic field of a plane monochromatic wave at the interface of two dielectrics. The validity of (47) and (48), which are FDTD analogues of the boundary conditions at the interface of two magnetics, is proven in §5.3.
@@
Beyond the specific results presented, this work underscores the continued relevance of rigorous accuracy analysis for classical discretizations, even in an era of advanced structure-preserving methods. By quantifying interface-induced errors in the Yee scheme and interpreting them through a transition-layer model, we provide tools that are immediately applicable to industrial simulation, educational instruction, and the validation of next-generation Maxwell solvers.
Data Availability Statement
The data that support the findings of this study are available upon reasonable request from the authors.
@@ Acknowledgments
The authors express their sincere gratitude to John B. Schneider, Associate Professor of Washington State University for useful discussion of FDTD dispersion effects and all anonymous reviewers for their valuable suggestions that improved this work.
The study was performed within the framework of a state assignment of the FRC Komi SC UB RAS (topic No. 125020501562-1), and also with financial support from the Ministry of Science and Higher Education of Russia within Agreement No. 075‑15‑2025‑455 dated 26.05.2025 during the conduct of NEXAFS studies.